OpenStudy (anonymous):

The aspect ratio of a 52-inch screen TV is 16:9. Find the horizontal and vertical dimensions of the screen. (Round your answers to the nearest inch.) horizontal dimension = vertical dimension =

OpenStudy (anonymous):

16+9 = 25 52/25=2.08 16*2.08=33.28 9*2.08=18.72 Check: 18.72+33.28 = 52 All OK Hor: 33.28 Vert: 18.72

OpenStudy (radar):

When solving this problem, you need to know the convention for TV screen size. A 52 in. TV screen refers to the "diagonal" dimension. I realize the problem is a math exercise, but google "TV Screen size" and you will get an answer, then do the math.

OpenStudy (radar):

http://www.displaywars.com/32-inch-16x9-vs-52-inch-16x9 "Google" will show that a 52 in. TV with an aspect ratio of 16:9 will have a width of 45.32 in (Hor) and a height of 25.49 in. (Vert). Now it would be interesting to determine if the math yields the same results.

OpenStudy (anonymous):

@radar the horizontal is correct, but the vertical wasn't.. interesting

OpenStudy (anonymous):

@radar the vertical was actually 25...

OpenStudy (radar):

The article stated that the vertical was 25.49 and if you worked it out and got 25, I would think that is close enough. Unfortunately I haven't gotten close to the width value of 45.32. Back to the drawing board. Thanks for solving it.

OpenStudy (radar):

O.K. I went and worked it out and I came up with a height (vertical) of 25.49 and a width of 45.32. Used pythagorean theorem and the aspect ratio (16y=9x) As you can see the results were pretty much aligned with the Google link.